OpenStudy (anonymous):

Estimate the length of the radius, if the area of a circle is 75 mm2.

OpenStudy (anonymous):

area is pir^2 so \[pir^2=75\] \[r^2=75/\pi\] \[r^2=23.87\] \[r=4.88mm\]

OpenStudy (anonymous):

pi r^2=75mm^@ r^2=75/3.142 r^2=23.87 r=under root 23.87 r=4.88 mm

OpenStudy (anonymous):

@sami-21 : how do you write in this style can u please tell me...:)

OpenStudy (anonymous):

@Qibtiya use the equation at the bottom of answer box

OpenStudy (anonymous):

@sami-21 i do but it does not turn out correct

OpenStudy (anonymous):

just matter of time you will get use to to it .